Circular No. 6170

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995N IN MCG –02-38-017 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova (V ∼ 17.5) on two Technical Pan Schmidt films taken on May 5.95 and 8.03 UT by D. Albanese and himself. SN 1995N is located at α = 14^{h}49^{m}28^{s}.27, δ = −10^{o}10′15′′.4 (equinox 2000.0), which is 34′′ west and 19′′ north of the brightest component of this irregular galaxy complex (also known as Arp 261). A nearby star of mag 17–18 is at position end figures 29^{s}.19, 37′′.3. S. Benetti, P. Bouchet, and H. Schwarz, European Southern Observatory (ESO), report that preliminary inspection of a partiallycalibrated CCD spectrogram (range 514.4–692.5 nm, resolution 0.8 nm), obtained on May 9.08 UT with the ESO 3.6-m telescope (+ EFOSC1), confirms that SN 1995N is a peculiar type-II supernova. Hα consists of one component-symmetric emission (peak measured at 660.1 nm; FWHM = 3.4 nm). Its profile looks similar to that shown by SN 1993N, ∼ 10 months after discovery. Narrow emission lines of Fe II at 519.8 nm; blend of [Fe II], [Fe III], and [Fe VII] at 530.6 nm; [Ca V] at 534.6 nm; [N II] at 579.2 nm; He I at 591.0 nm; [Fe VII] at 612.6 nm; [O I] at 634.1 nm; and [Fe X] at 641.5 nm are also present in the supernova spectrum. The last two lines are superimposed on a broad bump (with peak at ∼ 634.8 nm; FWHM ∼ 9.0 nm), probably associated with fast-moving oxygen from the supernova envelope. The coronal lines detected in this spectrum were also observed in SN 1988Z (Turatto et al. 1993, MNRAS 262, 128). The redshift of the parent galaxy (deduced from the H II-region lines present in the galaxy spectrum) is 0.0060.

ALEXIS J1644–032 D. Roussel-Dupré, on behalf of the ALEXIS team, reports: “Early on

May 4 UT, the ALEXIS extreme-ultraviolet satellite observed the beginning of a transient event. The location of the transient is α = 16^{h}44^{m}54^{s}, δ = −3^{o}12′.5 (equinox 2000.0; position accuracy ± 0^{o}.5). The transient was observable in the narrow, 66-eV channel (18.3 nm) but not the narrow, 93eV channel (13.0 nm); it initially appeared about as bright in Telescope 2B as HZ 43, which was being observed simultaneously and remained at that level throughout May 5. Both sources had a countrate of ∼ 0.15 photon cm^{−2} s^{−1} A^{−1}. The new source is currently near the edge of the field of view, making precise flux and position measurements difficult. Few good candidates for optical counterparts exist within the error box. Follow-up observations are encouraged.”

1995 May 10 Daniel W. E. Green